Book Publishing:

You are the owner of a publishing firm and you have a new author that you plan to publish. It is an action/espionage novel. You believe that the author has a good book, but it is her first book and you don’t really know what the sales numbers will look like. As such, you want to do a break even analysis to find out how many books you have to sell in order to get back your initial investment.

The book will be published in paperback sized 6”x9”. The initial set up cost for setting up the press for the book is $900 dollars. After the additional cost for setting up the press, each book will cost $2.19 each to make.

a) How much will it cost to print _ books?

o 100 books

o 200 books

o 500 books

o 1000 books

o 2000 books

b) What is the cost per book for each quantity
Answer #1

a)
for 100 books
cost = 900 + 2.19*100 = 1119

cost per book = 1119/100 = 11.19

for 200 books
cost = 900 + 2.19*200 = 1338

cost per book = 1338/200 = 6.69

for 500 books
cost = 900 + 2.19*500 = 1995

cost per book = 1995/500 = 3.99

for 1000 books
cost = 900 + 2.19*1000 = 3090

cost per book = 3090/1000 = 3.09

for 2000 books
cost = 900 + 2.19*2000 = 5280

cost per book = 5280/2000 = 2.64